OpenStudy (anonymous):

solve equation 2a² - 2a =1

OpenStudy (ranga):

Bring the constant term on the right to the left. Then use the quadratic formula to solve for a. (Note: the variable "a" used in this equation should not be confused with the "a" term used in the quadratic formula).

OpenStudy (anonymous):

confused????

OpenStudy (ranga):

2a^2 - 2a = 1 add -1 to both sides: 2a^2 - 2a - 1 = 0 Use the quadratic formula \[\Large \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]to find the roots.

OpenStudy (ranga):

\[\Large \frac{ -(-2) \pm \sqrt{(-2) ^{2} - 4(2)(-1)} }{ 2(2) } = \frac{ 2 \pm \sqrt{4 + 8} }{ 4 }\] = \[\Large \frac{ 2 \pm \sqrt{12} }{ 4 } = \frac{ 2 \pm \sqrt{4*3} }{ 4 } = \frac{ 2 \pm 2*\sqrt{3} }{ 4 } = \frac{ 1 \pm \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

thank you you save my life

OpenStudy (ranga):

The two roots are: (1 + sqrt(3)) / 2 and (1 - sqrt(3)) / 2 You are welcome.